# A Sanity Check for AI-generated Image Detection

### v2

@@ 5.4 Ablation Studies
Patchwise Feature Extraction. As shown in Table 7, removing either the high-frequency or the low-frequency patches results in obvious performance degradation in terms of accuracy. Without the high-frequency patches, the proposed method is unable to discern that the high-frequency regions of AI-generated images are smoother than those of real images, resulting in performance degradation. Similarly, without the low-frequency patches, the method cannot extract the underlying noise information, which is crucial for identifying AI-generated images with higher fidelity, leading to incorrect predictions.
@@
Semantic Feature Extraction. As shown in Table 7, the performance degrades significantly (76.70% vs 92.77%) when we remove the semantic branch. Intuitively, if the branch for semantic information extraction is absent, our method struggles to effectively capture images with semantic artifacts, resulting in significant drops.
@@
Visualization. To vividly demonstrate the effectiveness of our modules patchwise feature extraction (PFE) and semantic feature extraction (SFE), we conducted a visualization, as depicted in Fig. 4. In first row, the absence of semantic feature extraction results in many images with evident semantic errors going undetected. Similarly, the second row shows that, without patchwise feature extraction, numerous images lacking semantic errors still contain differing underlying information that remains unrecognized. Overall, our method, AIDE, achieves the best performance.
@@ 6 Conclusion
Potential societal impacts. Given that Chameleon demonstrates the capability to surpass the ”Turing Test”, there exists a significant risk of exploitation by malicious entities who may utilize AI-generated imagery to engineer fictitious social media profiles and propagate misinformation. To mitigate it, we will require all users of Chameleon to enter into an End-User License Agreement (EULA). Access to the dataset will be contingent upon a thorough review and subsequent approval of the signed agreement, thereby ensuring compliance with established ethical usage protocols.

### v3

@@ 5.4 Ablation Studies
Patchwise Feature Extraction. As shown in Table 5.2, removing either the high-frequency or the low-frequency patches results in obvious performance degradation in terms of accuracy. Without the high-frequency patches, the proposed method is unable to discern that the high-frequency regions of AI-generated images are smoother than those of real images, resulting in performance degradation. Similarly, without the low-frequency patches, the method cannot extract the underlying noise information, which is crucial for identifying AI-generated images with higher fidelity, leading to incorrect predictions.
@@
Semantic Feature Extraction. The performance degrades significantly (76.70% vs 92.77%) when we remove the semantic branch, as shown in Table 5.2. Intuitively, if the branch for semantic information extraction is absent, our method struggles to effectively capture images with semantic artifacts, resulting in significant drops.
@@
Visualization. To vividly demonstrate the effectiveness of our modules patchwise feature extraction (PFE) and semantic feature extraction (SFE), we conducted a visualization, as depicted in Fig. 9. In the first row, the absence of semantic feature extraction results in many images with evident semantic errors going undetected. Similarly, the second row shows that, without patchwise feature extraction, numerous images lacking semantic errors still contain differing underlying information that remains unrecognized. Overall, our method, AIDE, achieves the best performance.
@@ 6 Conclusion
Acknowledgements
WX would like to acknowledge the National Key R&D Program of China (No. 2022ZD0161400).
Appendix
Appendix A Experimental Details
A.1 Detectors
We choose a set of representative methods in AI-generated detection as baselines for comparison, including frequency-based (Frank et al., 2020; Liu et al., 2022a; Zhong et al., 2023), gradient-based (Tan et al., 2023), semantic-based (Ojha et al., 2023), reconstruction-based (Wang et al., 2023a), etc.
  • CNNSpot (CVPR’2020) (Wang et al., 2020) proposes that a naïve image classifier with simple data augmentations (i.e., JPEG compression and Gaussian blur) can generalize surprisingly to images generated by unknown GAN-based architectures.
  • FreDect (ICML’2020) (Frank et al., 2020) observes significant artifacts in the frequency domain of GAN-generated images and makes use of these artifacts for classification.
  • Fusing (ICIP’2022) (Ju et al., 2022) designs a two-branch model to fuse global spatial information and local informative features for training the classifier.
  • LNP (ECCV’2022) (Liu et al., 2022a) proposes to extract the noise pattern of images with a learnable denoising network and uses noise patterns to train a classifier.
  • LGrad (CVPR’2023) (Tan et al., 2023) employs gradients computed by a pretrained CNN model to present the generalized artifacts for classification.
  • UnivFD (CVPR’2023) (Ojha et al., 2023) uses CLIP features to train a binary liner classifier.
  • DIRE (ICCV’2023) (Wang et al., 2023a) observes obvious differences in discrepancies between images and their reconstruction by DMs and uses this feature to train a classifier.
  • PatchCraft (Arxiv’2024) (Zhong et al., 2023) compares rich-texture and poor-texture patches from images and extracts the inter-pixel correlation discrepancy as a universal fingerprint for classification.
  • NPR (CVPR’2024) (Tan et al., 2024) contributes to the architectures of CNN-based generators and demonstrates that the up-sampling operator can generate generalized forgery artifacts that extend beyond mere frequency-based artifacts.
Figure 3: Hyperparameter ablation of patch size and patch number introduced in our method.
A.2 Statistics of Public Benchmarks
Table 10 and Table 11 provide a detailed explanation of Benchmark AIGCDetectBenchmark (Zhong et al., 2023) & GenImage (Zhu et al., 2024) introduced in our main paper. There are two main benchmarks here: AIGCDetectBenchmark and GenImage. AIGCDetectBenchmark: It is trained on ProGAN and then tested on 16 different test sets, including data generated by both GAN and Stable Diffusion models. GenImage: It is trained on Stable Diffusion V1.4 and tested on a large amount of data generated by Stable Diffusion, with only a small portion of GAN data included. The test sets related to Stable Diffusion in AIGCDetectBenchmark are consistent with those used in GenImage.
Table 10: Statistics of Benchmark AIGCDetctBenchmark. SD and WFIR refer to Stable Diffusion and whichfaceisreal, respectively. The term ”Number” only counts on fake images and an equal number of real images is added for each generative model from the same source.
Generator Image Size Number Source
Train ProGAN (Karras et al., 2018) $256\times 256$ 360.0k LSUN (Yu et al., 2015)
 Test ProGAN (Karras et al., 2018) $256\times 256$ 8.0k LSUN (Yu et al., 2015)
StyleGAN (Karras et al., 2019) $256\times 256$ 12.0k LSUN (Yu et al., 2015)
BigGAN (Brock et al., 2019) $256\times 256$ 4.0k ImageNet (Deng et al., 2009)
CycleGAN (Zhu et al., 2017) $256\times 256$ 2.6k ImageNet (Deng et al., 2009)
StarGAN (Choi et al., 2018) $256\times 256$ 4.0k CelebA (Liu et al., 2015)
GauGAN (Park et al., 2019) $256\times 256$ 10.0k COCO (Lin et al., 2014)
StyleGAN2 (Karras et al., 2020) $256\times 256$ 15.9k LSUN (Yu et al., 2015)
WFIR (WFI, 2023) $1024\times 1024$ 2.0k FFHQ (Karras et al., 2019)
ADM (Dhariwal & Nichol, 2021) $256\times 256$ 12.0k ImageNet (Deng et al., 2009)
Glide (Nichol et al., 2022) $256\times 256$ 12.0k ImageNet (Deng et al., 2009)
Midjourney (mid, ) $1024\times 1024$ 12.0k ImageNet (Deng et al., 2009)
SD v1.4 (Sta, 2022) $512\times 512$ 12.0k ImageNet (Deng et al., 2009)
SD v1.5 (Sta, 2022) $512\times 512$ 16.0k ImageNet (Deng et al., 2009)
VQDM (Gu et al., 2022) $256\times 256$ 12.0k ImageNet (Deng et al., 2009)
Wukong (wuk, 2023) $512\times 512$ 12.0k ImageNet (Deng et al., 2009)
DALLE 2 (Ramesh et al., 2022) $256\times 256$ 2.0k ImageNet (Deng et al., 2009)
Table 11: Statistics of Benchmark GenImage. SD refers to Stable Diffusion. The term ”Number” only counts on fake images and an equal number of real images is added for each generative model from the same source.
Generator Image Size Number Source
Train SD v1.4 (Sta, 2022) $512\times 512$ 324.0k ImageNet Deng et al. (2009)
 Test BigGAN (Brock et al., 2019) $256\times 256$ 12.0k ImageNet (Deng et al., 2009)
ADM (Dhariwal & Nichol, 2021) $256\times 256$ 12.0k ImageNet (Deng et al., 2009)
Glide (Nichol et al., 2022) $256\times 256$ 12.0k ImageNet (Deng et al., 2009)
Midjourney (mid, ) $1024\times 1024$ 12.0k ImageNet (Deng et al., 2009)
SD v1.4 (Sta, 2022) $512\times 512$ 12.0k ImageNet (Deng et al., 2009)
SD v1.5 (Sta, 2022) $512\times 512$ 16.0k ImageNet (Deng et al., 2009)
VQDM (Gu et al., 2022) $256\times 256$ 12.0k ImageNet (Deng et al., 2009)
Wukong (wuk, 2023) $512\times 512$ 12.0k ImageNet (Deng et al., 2009)
Table 12: Comparison on the AIGCDetectBenchmark (Zhong et al., 2023) benchmark. Average precision (AP %) of different detectors (rows) in detecting real and fake images from different generators (columns). The best result and the second-best result are marked in bold and underline, respectively.
Method
ProGAN
StyleGAN
BigGAN
CycleGAN
StarGAN
GauGAN
StyleGAN2
WFIR
ADM
Glide
Midjourney
SD v1.4
SD v1.5
VQDM
Wukong
DALLE2
Mean
CNNSpot 100.00 99.83 85.99 94.94 99.04 90.82 99.48 99.85 75.67 72.28 66.24 61.20 61.56 68.83 57.34 53.51 80.41
FreDect 99.99 88.98 93.62 84.78 99.49 82.84 82.54 55.85 61.77 52.92 46.09 37.83 37.76 85.10 39.58 38.20 67.96
Fusing 100.00 99.50 90.70 95.50 99.80 88.30 99.60 93.30 94.10 77.50 70.00 65.40 65.70 75.60 64.60 68.12 84.23
LNP 99.89 98.60 84.32 92.83 100.00 78.85 99.59 91.45 94.20 88.86 76.86 94.31 93.92 87.35 92.38 96.14 91.85
LGrad 100.00 98.31 92.93 95.01 100.00 95.43 97.89 57.99 72.95 80.42 71.86 62.37 62.85 77.47 62.48 82.55 81.91
UnivFD 99.08 91.74 75.25 80.56 99.34 72.15 88.29 60.13 85.84 78.35 61.86 49.87 49.52 54.57 55.38 74.48 73.53
DIRE-G 58.79 56.68 46.91 50.03 40.64 47.34 58.03 59.02 99.79 99.54 97.32 98.61 98.83 98.98 98.37 99.71 75.54
DIRE-D 100.00 97.56 99.27 99.80 99.37 99.98 97.90 96.73 86.81 83.81 74.00 86.14 85.84 96.53 91.07 63.04 91.12
PatchCraft 100.00 98.96 99.42 85.26 100.00 81.33 97.74 95.26 93.40 94.04 96.48 99.06 99.06 96.26 97.54 99.56 95.84
NPR 100.00 99.81 87.87 98.55 99.90 85.57 99.90 65.38 74.61 85.73 85.41 84.02 84.67 81.20 80.51 76.72 86.87
AIDE 100.00 99.99 94.44 99.89 99.99 97.69 99.96 99.27 98.77 98.94 88.13 98.26 98.20 99.27 98.62 99.41 98.18
Appendix B More Experimental Results
B.1 AP Result
We additionally provide classification results regarding AP in Table 12. It is important to highlight that the AP (Average Precision) metric emphasizes different aspects compared to Acc (Accuracy). While Acc focuses on the overall correctness of predictions across all samples, AP provides a more comprehensive evaluation of a model’s performance across various thresholds, particularly in handling imbalanced datasets. On top of that, our method still achieves SOTA performance among these baselines on AP metric, which underscores the superiority of our approach. This indicates that our method not only excels in general prediction accuracy but also maintains robust performance across different decision thresholds, demonstrating its effectiveness in distinguishing between classes even in challenging scenarios.
B.2 More Ablation Studies
B.2.1 Patch Number and Patch Size
We further ablate some key parameters defined in our method, namely the size of patches (Patch Size) and the num of selected patches (Patch Number). As shown in Fig. 3 (a), both an excessively large and an overly small patch size can have certain impacts. If the patch size is too large, it may introduce additional irrelevant information, leading to interference with accurate judgment. On the other hand, if the patch size is too small, there may not be enough information available to make a proper judgment. As shown in Fig. 3 (b), for the patch number, the conclusion is that there is a correlation between the number of patches and the patch size.
B.2.2 ConvNeXt and ViT
We further explored the impact of CNN-Supervised network architectures such as CLIP-ConvNeXt and ViT-Supervised CLIP-ViT. The results of CLIP-ViT on Benchmark 1 have an average accuracy of 80.87%, which is significantly lower than the 92.77% achieved by CLIP-ConvNeXt. We speculate that this could be due to the CNN-supervised network architecture learning more low-level information, and for AI-generated image detection, low-level information is most crucial when the images are highly realistic.
Figure 4: Visualization of the effectiveness of PFE and SFE Modules with Grad-CAM (Selvaraju et al., 2017).
Appendix C Limitations
Although our method achieves state-of-the-art results on publicly available datasets and demonstrates highly competitive performance on Chameleon, its performance on our own dataset is still unsatisfactory. It leaves significant room for future improvement. Additionally, while the quality of our dataset is exceptionally high, its scale remains limited, and we aim to further expand its scale to better facilitate advancements in this field.
Appendix D Visualization
To more effectively verify the efficacy of the patchwise feature extraction (PFE) and semantic feature extraction (SFE) modules, we employ Grad-CAM (Selvaraju et al., 2017) to visualize the feature areas targeted by these modules. In Figure 4 (a), it is evident that the region highlighted by the red box exhibits distinct semantic issues, which our SFE module successfully captures with clarity. Conversely, in Figure 4 (b), there are no apparent semantic errors, and our PFE module accurately detects the low-level underlying noise information. Overall, our model, AIDE, demonstrates outstanding performance as a detector.
Appendix E Potential societal impacts
@@
Given that Chameleon demonstrates the capability to surpass the “Turing Test”, there exists a significant risk of exploitation by malicious entities who may utilize AI-generated imagery to engineer fictitious social media profiles and propagate misinformation. To mitigate it, we will require all users of Chameleon to enter into an End-User License Agreement (EULA). Access to the dataset will be contingent upon a thorough review and subsequent approval of the signed agreement, thereby ensuring compliance with established ethical usage protocols.